OpenStudy (anonymous):

Which numbers should be multiplied to obtain 175^2 - 124^2? 51 and 299 51 and 51 2601 and 2601 30501 and 30749 so i multiplied 175 by 175 and got 30,625 and then multiplied 124 by 124 and got 15,376 so im really confused bc none of the answers have to do with the numbers i got. can someone please help me?

hartnn (hartnn):

use this relation \(\Large a^2-b^2 = (a+b)(a-b)\)

hartnn (hartnn):

a= 175, b= 124 you just need a+b and a-b

OpenStudy (anonymous):



OpenStudy (anonymous):

so it would be A thank you so much!!

hartnn (hartnn):

yep, thats correct :) welcome ^_^

OpenStudy (anonymous):

wait i have one more question

OpenStudy (anonymous):

Which of the following is an actual zero of f(x) = x3 + 5x2 - 4x - 20? -20 -5 1 4 ok so i worked it out and got X=2, -5,-2 and i put the answe as -5 bc that was the only one there is that correct?

hartnn (hartnn):

yes, that is also correct :)

OpenStudy (anonymous):

ok thank you so much

hartnn (hartnn):

:)